OpenStudy (i0iz0gangster):

The results of an experiment are listed in several numerical forms listed below. Order the numbers from least to greatest. (3pts) 5^-3, 4/7 , squere root 5, 3/8 , 0.003

OpenStudy (lncognlto):

Are you allowed to use a calculator?

OpenStudy (i0iz0gangster):

i am afried not

OpenStudy (lncognlto):

Oh dear, okay then... here goes

OpenStudy (lncognlto):

5^-3 = 1/5^3 = 1/125 4/7 = 4/7 sqr 5 is, well greater than 2, so lets see 3/8 = 0.375 0.0003 =0.003

OpenStudy (i0iz0gangster):

a bit confused i must say if it not to much to ask mind clearing it up

OpenStudy (lncognlto):

I am just getting them into a comparable form. So then, in order of least to greatest, 0.003, 1/125, 0.375, 4/7, and finally sqr 2.

OpenStudy (lncognlto):

*sqr 5, I meant

OpenStudy (i0iz0gangster):

i thank dear sir/madam your help was much appretiotued and i shall reward you with a medal

OpenStudy (lncognlto):

Well, how kind of you to bestow such rewards upon me :)

OpenStudy (i0iz0gangster):

why its the least i can do when its so hard to find people like you in now a day society

OpenStudy (i0iz0gangster):

mind helping me with another one

OpenStudy (lncognlto):

Not at all...

OpenStudy (i0iz0gangster):

A pilot flies a plane south and then 600 miles west where she land the plane. How far south did the pilot fly the plane if she lands 610 miles from her starting point. and thank you for your kindness

OpenStudy (lncognlto):

Okay, this is a case for Pythagoras' theory. 610^2 (the hypotenuse of the triangle) = 600^2 + x^2 therefore x^2 = 372100 - 360000 =12100 tf x = sqr 12100 = 110 So she flew 110 miles south.

OpenStudy (i0iz0gangster):

again thank you dear sir your kindness is simply unmatched

OpenStudy (lncognlto):

Now, unfortunately, I must be one my way, as I have enthalpy changes which require my immediate assistance in solving themselves, but it has been a pleasure assisting such a kind person in distress.

OpenStudy (i0iz0gangster):

the pleasure is all mine dear sir

OpenStudy (lncognlto):

Nay, nay, but it is mine too.